The demand for Kia cars in a certain market is given by D(p) = 30 − p, MC = 5 (all prices and costs are expressed in units of thousands of dollars). Kia sells the cars to dealers at a wholesale price of w.

a If there is a single dealer with an exclusive right to retail Kia’s cars, find the profit-maximizing retail price it will charge, along with the quantity and the dealer’s profit-maximizing wholesale price w per car. What are the profit levels of both the manufacturer and the dealer

b Kia considers buying the dealer. What would the equilibrium price and profits be if it did both manufacture and retailing of its cars

c Define the double marginalization phenomenon. Illustrate this phenomenon graphically using your answers in parts (a) and (b)

d If, instead, there was perfect competition in retailing of Kia’s, at which level would the manufacturer set its wholesale price

e Compute Kia’s profits at the price in (d) and compare them against profits found in parts (a) and (b)


1
Expert's answer
2021-05-30T14:16:59-0400

a. If there is a single dealer, then the profit-maximizing retail price it will charge, along with the quantity are:

MR = MC,

p = 30 - Q,

MR = TR'(Q) = 30 - 2Q,

30 - 2Q = 5,

Q = 12.5 units,

p = 30 - 12.5 = 17.5.

b. If Kia considers buying the dealer, then the equilibrium price and profits will be higher if it did both manufacture and retailing of its cars.

c. Double Marginalization is the phenomenon in which different firms in the same industry that have their respective market powers but at different vertical levels in the supply chain (example, upstream and downstream) apply their own markups in prices.

d. If, instead, there was perfect competition in retailing of Kia’s, then:

MC = D,

30 - Q = 5,

Q = 25 units,

p = 30 - 25 = 5.

e. Kia’s profits at the price in (d) will be higher than profits found in parts (a) and (b).